a solution contains 6 % salt. How much water should be added to 90 ounces of the solution to make a 5.4% solution?

Answer by checkley79(3341)

.06*90=.054(90+x)
5.4=4.86+.054x
.054x=5.4-4.86
.054x=.54
x=.54/.054
x=10 oz. of water is needed.
Proof:
.06*90=.054(90+10)
5.4=.054*100
5.4=5.4